A bullet with a kinetic energy of 400 J strikes a wooden block where the bullet has a mass of 0.01kg. How fast was the bullet moving when it hits the wooden block
1
Expert's answer
2021-08-02T08:41:33-0400

The kinetic energy

"E_k=\frac{mv^2}{2}"

Hence, the speed of the bullet

"v=\sqrt{2E_k/m}=\sqrt{2*400\:\rm J/0.01\:\rm kg}=283\:\rm m/s"